If the mean height of bonsai trees is 52 cm with a standard deviation of 10 cm, 95% of the trees are between ______ and _______.


-115
-70
-65
-125
-85
-130

Answer by Theo(13342)

critical z-score is plus or minus 1.96.
z = (x-m)/s
z is the z-score
x is the raw score
m is the mean
s is the standad deviation
when z = 1.96 and s = 10 and m = 52, formula becomes 1.96 = (x-52)/10.
solve for x to get x = 1.96 * 10 + 52 = 71.6.
when z = -1.96 and s = 10 and m = 52, formula becomes -1.96 = (x-52)/10.
solve for x to get x = -1.96 * 10 + 52 = 32.4.
critical x-score is 52 plus or minus 19.6 = between 32.4 and 71.6
i don't see any of your selections matching that.